OpenStudy (anonymous):

is 3.87655 a rational number true or false

OpenStudy (anonymous):

can 3.87655 be put in fraction form?

OpenStudy (anonymous):

hmmm i do not know

OpenStudy (anonymous):

no but reoccurring ones are.

OpenStudy (anonymous):

if it ends isnt it rational?

OpenStudy (anonymous):

i think that if it ends on that 5, then it is rational but your calculator probably rounded to find that number

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

if you cant put it as a fraction it isnt rational

OpenStudy (anonymous):

thats wrong 2/4

OpenStudy (anonymous):

4/1 is a fraction too lol

OpenStudy (anonymous):

so its rational

OpenStudy (anonymous):

whats wrong?

OpenStudy (anonymous):

oh siiike

OpenStudy (anonymous):

you are just killin me today i thought you said if you can put it as a fraction

zepdrix (zepdrix):

\[\Large\rm 3.87655=\frac{3.87655}{1}\]Multiply the top and bottom of this fraction by 100,000. This will effectively move the decimal 5 places in the numerator.\[\Large\rm =\frac{387,655}{100,000}\] A rational number is one which can be written as the ratio of `integers`. Example:\[\Large\rm \frac{5}{2}\]5 over 2 is rational because 5 and 2 are integers. So for our problem, see how we were able to write it in that form? integer over integer.